OpenStudy (anonymous):

What type of triangle has side lengths of 4, 7, and 9?

OpenStudy (anonymous):

@maddierene13 I will be helping you today! Do you mean like acute, obtuse, right. Those kind of triangles? If you can elaborate more please ;)

OpenStudy (anonymous):

@hero Oh boy, its hero! ruler of open study

OpenStudy (anonymous):

use cosine rule.

hero (hero):

Use law of cosines to help you out with this \[c^2 = a^2 + b^2 - 2ab \cos C\]